you water the plants at a constant rate. after 5 seconds you have 58 oz in the watering can. 15 seconds later you have 28 oz of water. write an equation that represents the amount y (in oz) of water in the can after x seconds
Answer by JoelSchwartz(130)

y=amount of oz in can originally.
x=number of oz of water every second comes out of the container.
y-5x=58
y=5x+58
y-15x=28
5x+58-15x=28
-10x=28-58
-10x=-30
x=3
y=73